0215 UT\xA0\xA0\xA0\xA0 Dome dark.\xA0 Mostly cloudy.\xA0 Beginning to take slitless pixel flats.
0238 UT\xA0\xA0\xA0\xA0 Beginning MOS flatfield calibration for EAGES_J1426 mask
0257 UT\xA0\xA0\xA0\xA0 Beginning MOS arc lamp calibration for EAGES_J1426 mask
0258 UT\xA0\xA0\xA0\xA0 TO sent a test preset to TCS.\xA0 Telescope moving during some of the arc lamps.\xA0 Need to junk these and repeat.\xA0 See data file log.
0305 UT\xA0\xA0\xA0\xA0 Restarting arc lamps for EAGES_J1426
0318 UT\xA0\xA0\xA0\xA0 Opening dome & sending test preset to telescope
0334 UT\xA0\xA0\xA0\xA0 Moving to pointing star for initial check.\xA0 Clouds have cleared enough to begin work.
0347 UT\xA0\xA0\xA0\xA0 Slewing to Feige 34 for standard spectra
0408 UT\xA0\xA0\xA0\xA0 Moving to SN 2007rt.\xA0 Patchy clouds near horizon + high cirrus
0418 UT\xA0\xA0\xA0\xA0 Beginning spectrum of SN 2007rt reference object.\xA0 Cloud thickness (opacity) increasing
0528 UT\xA0\xA0\xA0\xA0 Moving to EAGES_J1426 field.\xA0 Continuing clouds of moderate thickness.\xA0 Unable to acquire R=14.5 guide star.
0637 UT\xA0\xA0\xA0\xA0 Trying again to acquire EAGES_J1426.\xA0 Guide star has returned.
0648 UT\xA0\xA0\xA0\xA0 Beginning science spectra of EAGES_J1426 mask.
0704 UT\xA0\xA0\xA0\xA0 Guide star stars going in and out.\xA0 Possible that cloud cover is too thick to allow good data.
0740 UT\xA0\xA0\xA0\xA0 Sky has largely cleared.\xA0 Second exposure on EAGES_J1426 should have high transparency.
0754 UT\xA0\xA0\xA0\xA0 Moving telescope to field for z5laeA.
0808 UT\xA0\xA0\xA0\xA0 Starting science integrations with z5laeA mask.\xA0 Guide star for this object is truncated at one side in WFS viewer; may experience collimation problems.
1013 UT\xA0\xA0\xA0\xA0 Losing collimation on guide star.\xA0\xA0 Stopping exposure early during mods1b.*.0038/mods1r.*.0047
1014 UT\xA0\xA0\xA0\xA0 Moving to EAGES_487588.
1016 UT\xA0\xA0\xA0\xA0 Aborting move and unwrapping telescope cables
1021 UT\xA0\xA0\xA0\xA0 Slew to EAGES_487588 failed when telescope went to wrong declination.\xA0 Re-trying.
1034 UT\xA0\xA0\xA0\xA0 Beginning science spectra on EAGES_487588
1141 UT\xA0\xA0\xA0\xA0 Finished EAGES_487588
1152 UT\xA0\xA0\xA0\xA0 Beginning bias frames & MOS mask calibration for z5laeA.

-- David Atlee - 13 Apr 2013
Topic revision: r6 - 13 Apr 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback